Motivation

Our Results/Contribution Summary

Title As It Is In the Proceedings

Include Only If Paper Has a Subtitle

F. Author^{1} S. Another^{2}

1Department of Computer Science University of Somewhere

2Department of Theoretical Philosophy University of Elsewhere

Conference on Fabulous Presentations, 2003

Motivation

Our Results/Contribution Summary

Outline

1 Motivation

The Basic Problem That We Studied Previous Work

2 Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Outline

1 Motivation

The Basic Problem That We Studied Previous Work

2 Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative. Use Uppercase Letters.

Subtitles are optional.

Use itemize a lot. Use very short sentences or short phrases.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

You can create overlays… using the pause command:

First item. Second item.

using overlay specifications:

First item. Second item.

using the general uncover command:

First item. Second item.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

You can create overlays… using the pause command:

First item. Second item.

using overlay specifications:

First item. Second item.

using the general uncover command:

First item. Second item.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

You can create overlays… using the pause command:

First item. Second item.

using overlay specifications:

First item. Second item.

using the general uncover command:

First item. Second item.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

You can create overlays… using the pause command:

First item. Second item.

using overlay specifications:

First item. Second item.

using the general uncover command:

First item. Second item.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

You can create overlays… using the pause command:

First item. Second item.

using overlay specifications:

First item. Second item.

using the general uncover command:

First item. Second item.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

You can create overlays… using the pause command:

First item. Second item.

using overlay specifications:

First item. Second item.

using the general uncover command:

First item. Second item.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Outline

1 Motivation

The Basic Problem That We Studied Previous Work

2 Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

Motivation

Our Results/Contribution Summary The Basic Problem That We Studied Previous Work

Make Titles Informative.

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Outline

1 Motivation

The Basic Problem That We Studied Previous Work

2 Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Make Titles Informative.

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Make Titles Informative.

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Make Titles Informative.

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Outline

1 Motivation

The Basic Problem That We Studied Previous Work

2 Our Results/Contribution

Main Results Basic Ideas for Proofs/Implementation

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Make Titles Informative.

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Make Titles Informative.

Motivation

Our Results/Contribution Summary Main Results Basic Ideas for Proofs/Implementation

Make Titles Informative.

Motivation

Our Results/Contribution

Summary

Summary

The first main message of your talk in one or two lines. The second main message of your talk in one or two lines. Perhaps a third message, but not more than that.

Outlook

Something you haven’t solved. Something else you haven’t solved.

Appendix  For Further Reading

For Further Reading I

A. Author. Handbook of Everything. Some Press, 1990.

S. Someone. On this and that. Journal of This and That, 2(1):50–100, 2000.